Solve each system by the addition method.
3x-2.5y=7.125
2.5x-3y=7.3125

Answer by stanbon(75887)

Solve each system by the addition method.
3x-2.5y=7.125
2.5x-3y=7.3125
---------------------
Multiply 1st by 2.5 and 2nd by -3 to get:
7.5x - 6.25y = 2.5*7.125
-7.5x +9y = -3*7.3125
---------------------------------------------
Add to get:
2.75y = -4.125
y = -1.5
---------------
Substitute to solve for "y":
3(-1.5) -2.5y = 7.125
-4.5 -2.5y = 7.125
-2.5y = 11.625
y = -4.65
======================
Cheers,
Stan H.
-------------------
Cheers,
Stan H.